OpenStudy (rissyroo118):

Simplify ; b^2c^5 b^2c^3 A)b^2 B)bc C)bc^2 D)c^2

Nnesha (nnesha):

\[\huge\rm \frac{ b^2 c^5}{ b^2 c^3 }\] like this ?

OpenStudy (rissyroo118):

exactly like that

Nnesha (nnesha):

hint! again EXPONENT RULES!! \[\huge\rm \frac{x^n }{ x^m} =x^{n-m}\] same base at the bottom and at the top so you can move denominator exponent to the top like in this example when you do that sign must change

Nnesha (nnesha):

exponent rules!! http://openstudy.com/study#/updates/54ff9b2ce4b064b5a94d524c very immportant

Nnesha (nnesha):

important*

OpenStudy (rissyroo118):

i wanna say it is A

OpenStudy (butterflydreamer):

That's not quite right LOL. It's not A :)

OpenStudy (butterflydreamer):